OpenStudy (anonymous):

Guys help needed how to solve limits involving two variables such as this ; lim (x,y->0,1) { ( x^2 +1) / ( x^2 + (y-1)^2 ) } Thanks in advance .

OpenStudy (anonymous):

...

OpenStudy (anonymous):

Pick any path and parameterize it

OpenStudy (anonymous):

Example :

OpenStudy (anonymous):

$$\lim_{(x,y)\to(0,1)}\frac{x^2+1}{x^2+(y-1)^2}=\lim_{y\to1}\frac1{(y-1)^2}=\infty$$Let approach along the \(y\)-axis

OpenStudy (anonymous):

Got that ; thanks the internet .

OpenStudy (anonymous):

Better http://tutorial.math.lamar.edu/Classes/CalcIII/Limits.aspx

OpenStudy (anonymous):

Yes... you pick any path and they should yield invariant limits

OpenStudy (anonymous):

Thank you.